The density of gas is 1.946 g dm-3 at 273 k and 76 cm Hg the gas is ?

Grade:9

1 Answers

Arun
5 years ago
Dear Afzal
 
Density=PM/RT==> M=dRT/P==>M=(1.964×10^(-3) × 82.1×273)/1==>M=44gTherefore, the gas is co2.76cmHg is approximately equal to 1 bar. Therefore, pressure taken is 1 bar.
 
Regards
Arun